Evaluate the Riemann sum for f(x)=6-2x find riemann sum under the area of curve with n=4 and equal subintervals [1,4] b)determine the actual under area under curve and explain why your answer in b underestimate
1
Expert's answer
2020-12-03T07:20:28-0500

We will approximate.

"\intop_1^4 (6-2x)\delta x"


Note that "f(x)=6-2x" and "a=1" and "b=4"

"n=4", so "\Delta x =\frac{b-a}{n}=\frac{4-1}{4}=\frac{3}{4}" or "0.75"


All endpoints: start with "a" and add "\u0394x" successively: "1,\ 1.75,\ 2.5,\ 3.25,\ 4"

The sub-intervals are;


"(1,1.75),\ (1.75, 2.5),\ (2.5, 3.25),\ (3.25, 4)"

we will use the left endpoint of each sub-interval.

The left endpoints are;


"1,\ 1.75,\ 2.5,\ 3.25"

Now the Riemann sum is the area of the 4 triangles. We find the area of the rectangle by


"Height \times Base = f(sample\ point) \times \Delta x"

Here we are using left endpoints of sub-intervals for sample points and "\Delta x=0.70", so


"R=(f(1)+f(1.75)+f(2.5)+f(3.25)+f(4)) \cdot 0.75"




"R=(4+2.5+1+(-0.5)+(-2))\times 0.75=3.75 units ^2"

b)determine the actual under area under curve and explain why your answer in b underestimate

The actual area under the curve is

"\intop _1^4 (6-2x)\delta x \implies [6x-x^2]_1^4\\
=[6(4)-(4)^2]-[6-1] \implies 8-5\\
\therefore Area= 3\ units^2"

The actual area is underestimated because I used the Left Riemann Sum and the function f(x) was decreasing on an interval